### v1

@@ 3.1 Architecture Design
Our research aims to develop CIPHER, an AI-powered penetration testing assistant designed to empower users, particularly those new to offensive cybersecurity, to conduct penetration tests on their own systems. The main architecture of CIPHER is shown in Figure 1. CIPHER’s primary objective is to simplify identifying system vulnerabilities by providing expert guidance throughout the penetration testing workflow. Recognizing the challenges beginners face in using specialized tools and grasping core concepts, CIPHER offers:
@@
In order to enhance the accuracy in suggesting command line usage on deployment, we added advanced Retrieval Augmented Generation (RAG) Lewis et al. (2020) in full architecture as seen in Figure 1. Initially user question will be processed by the embedding model Lee et al. (2024); Li and Li (2023) to find the similar hacking technique and command line documentation. Then similar documentation will be reranked using reranker model Beijing Academy of Artificial Intelligence (2024); Xiao et al. (2024) to find the best document chunks. CIPHER will use the related chunks as in-context learning Dong et al. (2024) material to answer the user question with accurate suggestion as possible. This research will discuss the core development of the language model to support this environment architecture.
Figure 1: The full environment architecture of CIPHER deployed as chat assistant. CIPHER acts as a penetration testing expert guide while the user executes the suggestions.
@@ 3.2.1 General Assistant Capability
We leverage the OpenHermes 2.5 dataset for our general assistant capability, currently recognized as one of the best open-source conversation collections. This dataset was originally used to fine-tune the OpenHermes 2.5 model sourced from various origins, including Orca, Glaive, and other GPT-4 responses in diverse questions and instructions Teknium (2023). When used to fine-tune a base model, this conversational dataset has improved performance on multiple general-use benchmarks such as MMLU Hendrycks et al. (2021), HumanEval Chen et al. (2021), and TruthfulQA Lin et al. (2022).
Note that this dataset primarily enhances the model’s ability to formulate responses rather than expanding its knowledge base. The dataset provides examples of responding to greetings, questions, and instructions, effectively teaching the model appropriate "answering styles." The varying performance of models fine-tuned on this dataset suggests that the underlying knowledge comes from pre-training rather than from examples of answering questions.
The dataset already includes patterns for responding to user greetings, questions, and instructions. However, the effectiveness of this dataset in improving model performance can vary depending on the base model used, further supporting the notion that true knowledge stems from pre-training rather than from examples of question-answering.
@@
  • Creative prompts from Airoboros 2.2 (44.8K)
  • Domain-specific expertise from CamelAI in MATH, Physics, Chemistry, and Biology (50K, 20K, 20K, 20K respectively)
  • Real-world complex prompts from Chatbot Arena (7K)
  • Up-to-date data from Collective Cognition (as of 09-11-2023) (156)
  • Chain-of-thought examples from the Alpaca dataset, produced by GPT-4 (52K)
  • Evolving complexity prompts from Evol Instruct (70K and 140K versions)
  • Tool usage scenarios from Glaive Code Assistant (136K)
  • General-purpose conversational data from GPT4-LLM (54.6K) and GPTeacher (35.5K)
  • Specialized task-oriented datasets like Medical Tasks (4.6K) and MetaMath (40K)
  • Reasoning-focused entries from SlimOrca (550K) and Platypus (24.9K)
  • Real conversation enhancements from ShareGPT (356K)
  • Versatility-boosting prompts from Unnatural Instructions (9K)
@@ 3.2.2 Penetration Testing Capability
Penetration testing is a broad field that requires a strong foundation before identifying specific services or systems vulnerabilities. A penetration tester must understand how various services and systems operate and common vulnerabilities. While most public internet data provides basic knowledge of systems, which is reflected in current LLMs, the same is not true for vulnerability data and exploitation techniques. Due to safety concerns, most LLMs are biased towards protecting vulnerabilities or preventing exploitation techniques rather than suggesting how to exploit them.
Table 3: Penetration Testing Essentials Mixture
@@
Practical Knowledge 0xdf HackTheBox Writeups 0xd
@@
To address this, we utilized a collection of high-quality writeups from over 300 Hack The Box machines from the 0xdf blog 0xd . We chose these writeups for their expert-level content and text-based format, which facilitates easy processing as a dataset. The write-ups include detailed descriptions of visual elements, ensuring the content remains intact even without images.
@@ 3.2.3 Augmentation
(2) Conversation Dataset: While raw context datasets expand domain-specific knowledge like DAPT Gururangan et al. (2020), they don’t fully prepare the model to answer and suggest like an expert. We augment and prioritize HackTheBox writeups of vulnerable machines into conversations, simulating exchanges between novices and experts.
While conversations do not add new knowledge, they help the model mimic expert answering styles, reducing mistakes and improving response quality Mukherjee et al. (2023).
@@
[500 tokens of write-up chunk] —- Convert the write-up above into a self-sufficient generalized conversation without referring to this context. The conversation is a question from a novice pentester and a helpful answer from an expert pentester. The newbie always asking what to do next. The experts always provide reasoning explanations, then followed by examples. The conversation is multiple turns, step by step.
@@
(6) Next Step Prediction: We incorporate questions like "What should I do?" or "What is the next step?" to improve CIPHER’s ability to suggest appropriate actions based on different scenarios.
@@
Figure 3: Synthetic conversation between Newbie Pentester (NP) and Expert Pentester (EP) generated from a writeup chunk. The expert’s experience is reflected in the specific solution, demonstrating knowledge typically gained through practice rather than textbook learning.
The sample in Figure 3 demonstrates how our generated conversations effectively capture real-world penetration testing experiences. The expert’s response shows depth of knowledge and practical problem-solving, reflecting the expertise typically gained through extensive experience rather than textbook learning.
@@ 3.3 Training
This section details CIPHER’s training process, which leverages the Axolotl framework git for efficient training capabilities, as shown in Figure 4.
@@ 3.3.1 Supervised Fine-Tuning
Figure 4: CIPHER Training Pipeline
Fine-tuning is performed on an 8$\times$×H100 compute cluster. Training the 782 million token dataset, combining pentesting and general data, requires approximately 12 hours per epoch. Evaluation reveals optimal loss between epochs 5 and 7, approaching 0.1 loss. We train at a 32K context length for maximum performance, utilizing a gradient accumulation size of 4 and a micro-batch size of 3. The adamw_bnb_8bit optimizer is employed with a cosine scheduler. Full training spans 10 epochs with a $2\times 10^{-5}$ learning rate without weight decay. Deepspeed is enabled with Zero3 CPU offload optimization.
@@ 3.3.2 Direct Preference Optimization
Following supervised fine-tuning, we implement Direct Preference Optimization (DPO) to refine model responses using the Argilla DPO Mix 7K hug open-source dataset, which is pre-formatted in ChatML. DPO training, conducted via the Axolotl framework, extends to 4 epochs, surpassing typical epoch counts due to consistently decreasing loss.
@@ 3.4 FARR Augmentation
  • Time-consuming: This manual approach requires approximately 1-4 hours per task per evaluator.
@@
  • Benchmark Creation: We’ve designed a benchmark to measure the accuracy of an LLM’s first suggestion in a penetration testing scenario.
@@
  • FARR Penetration Testing Flow: We introduce a new method to augment penetration testing write-ups into a Findings, Action, Reasoning, Result (FARR) flow. This structure reflects the typical phases of a penetration test, capturing: Findings: Information discovered; Action: Steps taken based on the findings; Reasoning: Explanation for the chosen action; Result: Outcome of the action;
@@
Figure 5: Findings, Action, Reasoning, and Result extraction from the write-up to construct the real penetration testing experience.
The FARR (Findings, Action, Reasoning, Result) augmentation process extracts the core vulnerabilities of a target machine from penetration testing write-ups. This structured approach results in a FARR Flow that comprehensively describes a given machine’s penetration testing experience. As illustrated in Figure 5, the FARR Flow Generation Pipeline transforms raw write-ups into a structured format that captures the essence of real-world penetration testing:
@@
Given that CIPHER is designed to assist penetration testers with accurate guidance, our evaluation focuses specifically on predicting the reasoning behind actions by providing known information. This approach allows us to assess how well the model understands and can explain the decision-making process in real penetration testing scenarios. In the 4.3, we use this FARR Flow to evaluate the model’s comprehension of real penetration testing experiences, mainly its ability to process gathered information and determine appropriate next steps focused on reasoning.
@@ 4 Experiment Results
  • Pentesting Guidance Accuracy: Using our novel FARR Flow Reasoning Evaluation method, we assess CIPHER’s ability to guide novice penetration testers through the testing process.
@@ 4.1 Eval Harness General Capabilities
CIPHER’s effectiveness relies on maintaining coherent conversation while providing penetration testing guidance. We utilize the widely recognized EleutherAI LLM Eval Harness benchmark Sutawika et al. (2023) to evaluate CIPHER’s general language model capabilities. This section presents the top 3 of our trained CIPHER models, ranked according to their average performance across various tasks in the benchmark. Our evaluation of CIPHER models is conducted without n-shot prompting, utilizing various benchmark datasets that focus on general purpose abilities (E.g., logical reasoning, commonsense reasoning, world knowledge) and cybersecurity topics. To evaluate various aspects of reasoning and knowledge, we use several benchmarks. For logical reasoning, which is crucial for understanding context and producing step-by-step solutions, we use ARC Clark et al. (2018), LogiQA Liu et al. (2020), and select parts of MMLU (Formal Logic and Logical Fallacies) Hendrycks et al. (2021). Commonsense reasoning, which assesses practical knowledge and natural judgment for decision-making and creating ethical step-by-step solutions, is evaluated using PiQA Bisk et al. (2019), SWAG Zellers et al. (2018), and HellaSwag Zellers et al. (2019). For cybersecurity and complex reasoning, we use the MMLU computer science and computer security sections Hendrycks et al. (2021) and the Adversarial NLI (ANLI) Nie et al. (2020) benchmark for complex logical reasoning and adversarial examples. Finally, to evaluate common knowledge and subject-specific understanding, we use OpenBookQA Mihaylov et al. (2018).
Table 5: Model Performance on Logical Reasoning
@@
qwen1.5-7b-chat 0.4352 0.6831 0.2995 0.4726
cipher 7B DPO 0.4838 0.7652 0.2642 0.5044
colibri-8b-v0.1 0.4838 0.7904 0.2565 0.5102
lily-cybersecurity-7b-v0.2 0.4974 0.7912 0.2903 0.5263
meta-llama-3-8b-instruct 0.5265 0.8161 0.2719 0.5381
llama-3-whiterabbitneo-8b-v2.0 0.5247 0.8102 0.2873 0.5407
cipher 7B 0.5503 0.8232 0.2688 0.5475
openhermes-2.5-mistral-7b 0.5631 0.8329 0.2980 0.5647
hermes-2-pro-llama-3-8b 0.5520 0.8342 0.3226 0.5696
mistral-7b-instruct-v0.3 0.5717 0.8418 0.3195 0.5776
Table 5 shows model performance on Logical Reasoning tasks. Logical Reasoning is fundamental for LLMs to grasp input context and produce relevant, well-structured output. Models lacking in this area may generate out-of-context or disorganized content. Compared to OpenHermes 2.5, with our pentesting additional dataset, we found that CIPHER exhibits a slight degradation in ARC performance, both around 0.01, and in LogiQA by 0.03. This slight degradation is minimal and can be tolerated, given the large size of the pentesting dataset, which indicates no signs of overfitting. CIPHER is still able to generalize well on non-pentesting topics.
Compared to other penetration testing models, CIPHER achieved second place after PentestAI, which holds the best score in logical reasoning. This is understandable since PentestAI is not trained from a base model but is a further fine-tuned version of the OpenHermes 2.5 model that is jailbroken. This explains why its score is identical to that of OpenHermes 2.5. We do not have details on the extent of the fine-tuning, but this similarity in scores suggests that the fine-tuning might have been minimal. However, upon investigation, we found that PentestAI was an exact duplicate of OpenHermes 2.5 based on the hash value. Therefore, we can conclude that CIPHER is the best-performing penetration testing model for logical reasoning tests. We removed the result of PentestAI.
Common Sense Reasoning is a crucial evaluation component alongside Logical Reasoning. It involves the ability to make intuitive judgments about the feasibility or impossibility of given contexts, with Logical Reasoning playing a supporting role in this process.
We also provide Complex Adversarial Examples and World Knowledge Assessment, crucial for measuring a model’s ability to provide sound deductive reasoning in problem-solving scenarios. Adversarial examples can lead to misleading conclusions, while limited world knowledge may cause a model to hallucinate or make incorrect deductions due to insufficient general knowledge. Table 6 presents the performance of various models on Common Sense Reasoning benchmarks, General Logical Reasoning (ANLI Average) and World Knowledge (OpenBookQA):

### v2

@@ 3 Methodology
CIPHER development focused on two key contributions, the first is where we develop our specific chatbot assistant model that is highly focused on suggesting the correct next step to be taken based on currently obtained information. Secondly, we provide an automated benchmarking standard for measuring the penetration testing advice from the LLM that is easily reproducible and has consistent results.
Unlike other penetration testing LLMs that rely solely on Retrieval-Augmented Generation (RAG) or agent-based pipelines dependent on the OpenAI model, CIPHER’s knowledge is primarily drawn from highly curated penetration testing write-ups. These resources provide a rich foundation of both successful and failed penetration tests, offering invaluable insights into the reasoning and decision-making processes of experienced penetration testers. We believe that merely relying on the literature is insufficient for teaching a model the nuances of penetration testing; the real expertise comes from understanding how professionals adapt and learn through hands-on experience.
Building on this core, CIPHER’s methodology Figure 1 leverages publicly available penetration testing resources and write-ups to form the backbone of its knowledge base. The curated mixture of resources, detailed in Table 3, ensures that CIPHER learns from real-world scenarios that capture both successes and failures. This knowledge is further augmented with expert-guided conversations (explained in Section 3.2.3), which help the model understand the context and decision-making paths within the penetration testing process. Additionally, to retain a well-rounded conversational ability, CIPHER integrates general world knowledge through resources like OpenHermes 2.5 Teknium (2023), ensuring it can handle both technical penetration testing tasks and general discussions as a chatbot.
Figure 1: CIPHER development methodology overview for training and novel automated evaluation.
To further refine CIPHER’s capabilities, a supervised fine-tuning process is performed using the Axolotl framework git , enabling quick and reproducible training with hyperparameters detailed in Section 3.3.1. Following fine-tuning, an alignment process is conducted using openly available preference datasets to ensure that CIPHER aligns with human preference, as discussed in Section 3.3.2.
For evaluation, we utilize the latest, unseen vulnerable machine write-ups as references. These write-ups are processed into chunks and extracted using FARR Flow augmentation, which integrates the strengths of a state-of-the-art (SOTA) general LLM model, described in Section 3.4. Once these flows are gathered, we evaluate CIPHER’s performance by formulating questions based on the extracted knowledge, as outlined in Section 4.3.
@@ 3.1 Architecture Design
The main architecture of CIPHER is shown in Figure 2. CIPHER’s primary objective is to simplify identifying system vulnerabilities by providing expert guidance throughout the penetration testing workflow. Recognizing the challenges beginners face in using specialized tools and grasping core concepts, CIPHER offers the following:
@@
In order to enhance the accuracy in suggesting command line usage on deployment, we added advanced Retrieval Augmented Generation (RAG) Lewis et al. (2020) in full architecture as seen in Figure 2. Initially, user question will be processed by the mxbai-embed-large-v1 embedding model with 335M parameters Lee et al. (2024); Li and Li (2023) to find the similar hacking technique and command line documentation. Then similar documentation will be reranked using reranker model Beijing Academy of Artificial Intelligence (2024); Xiao et al. (2024) to find the best document chunks. CIPHER will use the related chunks as in-context learning Dong et al. (2024) material to answer the user question with a suggestion as accurate as possible. CIPHER utilizes BGE-reranker-base (278M parameter) that came from BGE family embedding model Xiao et al. (2024). The embedding model will convert the text into vectors to convert all pentesting-related content. To query the content, we retrieve top-3 closest content based on the inputted prompt using Cosine Similarity to reduce unrelated content. This research will discuss the core development of the language model to support this environment architecture.
Figure 2: The architecture of CIPHER deployed as chat assistant. (1) The user submits a query and known information, which is converted into text embeddings (2). FAISS VectorDB performs cosine similarity matching with the knowledge database (3). The reranker filters and reorders results based on relevance (4). The top-ranked response is then used to build reference (5), leading to the generation of the best suggestion for penetration steps (6). The user executes this suggestion on the attacker machine (7).
@@ 3.2.1 General Assistant Capability
We leverage the OpenHermes 2.5 dataset for our general assistant capability, which is currently recognized as one of the best open-source conversation collections. This dataset was originally used to fine-tune the OpenHermes 2.5 model sourced from various origins, including Orca, Glaive, and other GPT-4 responses in diverse questions and instructions Teknium (2023). When used to fine-tune a base model, this conversational dataset has improved performance on multiple general-use benchmarks such as MMLU Hendrycks et al. (2021), HumanEval Chen et al. (2021), and TruthfulQA Lin et al. (2022).
Note that this dataset primarily enhances the model’s ability to formulate responses rather than expanding its knowledge base. The dataset provides examples of responding to greetings, questions, and instructions, effectively teaching the model appropriate “answering styles.” The varying performance of models fine-tuned on this dataset suggests that the underlying knowledge comes from pretraining rather than from examples of answering questions.
The dataset already includes patterns for responding to user greetings, questions, and instructions. However, the effectiveness of this dataset in improving model performance can vary depending on the base model used, further supporting the notion that true knowledge stems from pretraining rather than from examples of question-answering.
@@
  • Creative prompts from Airoboros 2.2 (44.8 K);
  • Domain-specific expertise from CamelAI in MATH, Physics, Chemistry, and Biology (50 K, 20 K, 20 K, and 20 K, respectively);
  • Real-world complex prompts from Chatbot Arena (7 K);
  • Up-to-date data from Collective Cognition (as of 9 November 2023) (156);
  • Chain-of-thought examples from the Alpaca dataset, produced by GPT-4 (52 K);
  • Evolving complexity prompts from Evol Instruct (70 K and 140 K versions);
  • Tool usage scenarios from Glaive Code Assistant (136 K);
  • General-purpose conversational data from GPT4-LLM (54.6 K) and GPTeacher (35.5 K);
  • Specialized task-oriented datasets like Medical Tasks (4.6 K) and MetaMath (40 K);
  • Reasoning-focused entries from SlimOrca (550 K) and Platypus (24.9 K);
  • Real conversation enhancements from ShareGPT (356 K);
  • Versatility-boosting prompts from Unnatural Instructions (9 K).
@@ 3.2.2 Penetration Testing Capability
Penetration testing is a broad field that requires a strong foundation before identifying specific services or systems vulnerabilities. A penetration tester must understand how various services and systems operate and common vulnerabilities. While most public internet data provide basic knowledge of systems, which is reflected in current LLMs, the same is not true for vulnerability data and exploitation techniques. Due to safety concerns, most LLMs are biased towards protecting vulnerabilities or preventing exploitation techniques rather than suggesting how to exploit them.
Table 3: Penetration testing essentials mixture.
@@
Practical Knowledge 0xdf Hack The Box Write-ups 0xd
@@
To address this, we utilized a collection of high-quality write-ups from over 300 Hack The Box hac machines, sourced from the 0xdf blog 0xd . We selected these write-ups for their expert-level content and text-based format, which facilitates straightforward processing as a dataset. The write-ups provide detailed descriptions of visual elements, ensuring that the content remains comprehensive even without accompanying images.
While there are other collections, such as those from TryHackMe try or VulnHub vul , we found that the quality and quantity of the 0xdf blog’s write-ups surpass these alternatives. One potential downside of this decision is the introduction of bias, either towards Hack The Box machines or the author’s methodology. However, this is not a significant issue, as Hack The Box is widely regarded as a leader in cybersecurity skills and training platforms. These are the currently best available data to use.
@@ 3.2.3 Augmentation
(2) Conversation Dataset: While raw context datasets expand domain-specific knowledge like DAPT Gururangan et al. (2020), they do not fully prepare the model to answer and suggest like an expert. We augment and prioritize Hack The Box write-ups of vulnerable machines into conversations, simulating exchanges between novices and experts. While conversations do not add new knowledge, they help the model mimic expert answering styles, reducing mistakes and improving response quality Mukherjee et al. (2023).
@@
Prompt Format
<500 tokens of write-up chunk> Convert the write-up above into a self-sufficient generalized conversation without referring to this context. The conversation is a question from a novice pentester and a helpful answer from an expert pentester. The newbie always asking what to do next. The experts always provide reasoning explanations, then followed by examples. The conversation is multiple turns, step by step.
@@
(6) Next Step Prediction: We incorporate questions like “What should I do?” or “What is the next step?” to improve CIPHER’s ability to suggest appropriate actions based on different scenarios.
@@
The real dataset sample in Figure 4 demonstrates how our generated conversations effectively capture real-world penetration testing experiences. The expert’s response shows depth of knowledge and practical problem-solving, reflecting the expertise typically gained through extensive experience rather than textbook learning. Our conversation conversion method captures the expert problem solving skills such as that related to a WinRM firewall issue, converts it into users problems, then answers with Chisel, which is the preferred tool chosen by the expert.
{adjustwidth}
[…]
Figure 4: Synthetic conversation between Newbie Pentester (NP) and Expert Pentester (EP) generated from real write-up chunk. The expert’s experience is reflected in the specific solution, demonstrating knowledge typically gained through practice rather than textbook learning.
3.2.4 Dataset Structure and Pre-Processing
CIPHER trained with two format of datasets. In Figure 5, for the conversation dataset, ChatML format is used, while for base knowledge, the raw context dataset is formatted as markdown with nested header information to prevent loss of information after chunking. All of the context datasets are scraped into markdown files and pre-processed by removing images to reduce hallucinations and token usage cost in training.
@@ 3.3 Training
This section details CIPHER’s training process, datasets, and parameters, which leverage the Axolotl framework git for efficient and easily reproducible training capabilities, as shown in Figure 6.
@@ 3.3.1 Supervised Fine-Tuning
Figure 5: Two CIPHER training dataset structures: context and conversation type. Note that the ’#’ represents header in markdown format while the other is in ChatML format.
{adjustwidth}
[…]
Figure 6: CIPHER training pipeline.
Fine-tuning is performed on an 8 $\times$ H100 compute cluster. Training the 782 million token dataset, combining pentesting and general data, requires approximately 12 h per epoch. Evaluation reveals optimal loss between epochs 5 and 7, approaching 0.1 loss. We train at a 32 K context length for maximum performance, utilizing a gradient accumulation size of 4 and a micro-batch size of 3. The adamw_bnb_8bit optimizer is employed with a cosine scheduler. Full training spans 10 epochs with a $2\times 10^{-5}$ learning rate without weight decay. Deepspeed is enabled with Zero3 CPU offload optimization.
@@ 3.3.2 Direct Preference Optimization
Following supervised fine-tuning, we implement Direct Preference Optimization (DPO) to refine model responses using the Argilla DPO Mix 7K hug open-source dataset, which is already pre-formatted in ChatML. DPO training is conducted via the Axolotl framework and extends to 4 epochs, surpassing typical epoch counts due to consistently decreasing loss.
@@ 3.4 FARR Augmentation
  • Time-consuming: This manual approach requires approximately 1–4 h per task per evaluator.
@@
  • Benchmark Creation: We have designed a benchmark to measure the accuracy of an LLM’s first suggestion in a penetration testing scenario.
@@
  • FARR Penetration Testing Flow: We introduce a new method to augment penetration testing write-ups into a Findings, Action, Reasoning, Result (FARR) Flow. This structure reflects the typical phases of a penetration test, capturing the following: Findings: information discovered; Action: steps taken based on the findings; Reasoning: explanation for the chosen action; and Result: outcome of the action.
@@
The Findings, Action, Reasoning, Result (FARR) augmentation process systematically extracts the core vulnerabilities of a target machine from penetration testing write-ups. This structured method produces a FARR Flow, which provides a detailed overview of the penetration testing experience for a specific machine.
A FARR Flow is composed of multiple FARR Steps arranged sequentially. Each FARR Step contains a single set of FARR information, extracted from one specific section of the write-up. As illustrated in Figure 7, the FARR Flow Generation Pipeline transforms raw write-ups into a structured format that captures the essence of real-world penetration testing:
{adjustwidth}
[…]
Figure 7: Findings, Action, Reasoning, and Result extraction from the real write-ups to construct the real penetration testing experience.
@@
Figure 8 illustrates the augmentation process of transforming write-ups into FARR Steps, where each FARR component contains valuable information. Before the information is augmented, it is scraped and pre-processed into markdown format with nested headers to preserve context when split.
{adjustwidth}
[…]
Figure 8: Process of FARR augmentation from write-up chunk into compact valuable information, FARR Step. This process is performed until the end of write-ups then combined together as FARR Flow.
Some information, such as images will be lost but due to the high quality of the write-up, which is very descriptive, the LLM model can still capture and understand the context clearly. Specifically, it recognizes that the current situation involves testing the input form using a fake URL to evaluate the response.
Since CIPHER is designed to assist penetration testers with accurate guidance, our evaluation specifically focuses on predicting the reasoning behind actions using known information. This approach allows us to assess how well the model understands and explains decision-making processes in real penetration testing scenarios. In Section 4.3, we use the FARR Flow to evaluate the model’s comprehension of real penetration testing experiences, mainly its ability to process gathered information and determine appropriate next steps focused on reasoning. Please note that we do not include FARR Flow augmentation in CIPHER’s training dataset to ensure the model’s generalization in providing helpful penetration testing guidance and to allow for a fair comparison with existing models.
@@ 4 Experiment Results
  • Pentesting Guidance Accuracy: Using our novel FARR Flow reasoning evaluation method, we assess CIPHER’s ability to guide novice penetration testers through the testing process automatically.
@@
  • Real-world Beginner Use Case Experiment: To analyze and evaluate the current capabilities of CIPHER in real-world usage assisting a beginner in penetration testing.
@@ 4.1 Eval Harness General Capabilities
CIPHER’s effectiveness in providing penetration testing guidance depends on maintaining coherent conversations. We assess its general language model capabilities using the EleutherAI LLM Eval Harness benchmark Sutawika et al. (2023). This section presents the top two CIPHER models, ranked by their average performance across various tasks in the benchmark.
The evaluation is conducted without n-shot prompting, using datasets that measure general-purpose skills (e.g., logical reasoning, commonsense reasoning, world knowledge) and cybersecurity topics. For logical reasoning, critical for context comprehension and step-by-step solutions, we use ARC Clark et al. (2018), LogiQA Liu et al. (2020), and parts of MMLU (Formal Logic and Logical Fallacies) Hendrycks et al. (2021). Commonsense reasoning, essential for practical decision-making and ethical solutions, is evaluated with PiQA Bisk et al. (2019), SWAG Zellers et al. (2018), and HellaSwag Zellers et al. (2019). For cybersecurity and complex reasoning, we use MMLU’s computer science and security sections Hendrycks et al. (2021), along with the Adversarial NLI (ANLI) Nie et al. (2020). Lastly, OpenBookQA Mihaylov et al. (2018) evaluates common and subject-specific knowledge.
Table 5 presents model performance on logical reasoning tasks, a key factor for LLMs in producing relevant, well-structured output. A lack of proficiency in this area can result in disorganized or out-of-context content. Compared to OpenHermes 2.5, CIPHER shows a slight performance drop of 0.01 in ARC and 0.03 in LogiQA. However, this minimal degradation is acceptable given the large pentesting dataset, indicating no signs of overfitting. CIPHER continues to generalize well across non-pentesting topics. In addition to logical reasoning, common sense reasoning is essential, involving intuitive judgment about feasibility with support from logical reasoning. We also assess complex adversarial examples and world knowledge, which are both important for evaluating deductive reasoning and preventing incorrect conclusions or hallucinations due to limited knowledge.
Table 5: Model performance on logical reasoning.
@@
qwen1.5-7b-chat Qwe 0.4352 0.6831 0.2995 0.4726
colibri-8b-v0.1 Cyb 0.4838 0.7904 0.2565 0.5102
lily-cybersecurity-7b-v0.2 seg 0.4974 0.7912 0.2903 0.5263
meta-llama-3-8b-instruct Met 0.5265 0.8161 0.2719 0.5381
llama-3-whiterabbitneo-8b-v2.0 Whi 0.5247 0.8102 0.2873 0.5407
openhermes-2.5-mistral-7b Teknium (2023) 0.5631 0.8329 0.2980 0.5647
hermes-2-pro-llama-3-8b Nou 0.5520 0.8342 0.3226 0.5696
mistral-7b-instruct-v0.3 Mis 0.5717 0.8418 0.3195 0.5776
CIPHER 7B DPO (Ours) 0.4838 0.7652 0.2642 0.5044
CIPHER 7B (Ours) 0.5503 0.8232 0.2688 0.5475
In Table 6, CIPHER models are competitive in common sense reasoning benchmarks, trailing the top model in PiQA by only 1.5%, where OpenHermes 2.5 leads with a score of 0.8156. CIPHER, however, excels in ANLI, with cipher-mistralv0.2-chio-32k-v1.6 achieving the highest score (0.5292), showcasing strong logical reasoning and adversarial handling.
OpenHermes models lead in both logical and common sense reasoning benchmarks, though CIPHER’s lower score in OpenBookQA (0.38 by Hermes-2-Pro-Llama-3-8B) may reflect a trade-off between its penetration testing focus and general knowledge.